# Computing the Topological Degree of Maps Between 2-Spheres

### v1

@@ Abstract.
We describe an effective method for computing the topological degree of continuous functions $R:S^{2}\to S^{2}$, where $S^{2}$ is the Riemann sphere. Our approach generalizes the degree formula for rational functions of complex polynomials, $\frac{f}{g}$, without common zeros. To apply our method, it is necessary to represent the function $R$ as the ratio of two continuous complex-valued functions $f$ and $g$ without common zeros. By utilizing Hopf fibration, this method reduces the problem to computing the winding number of a loop. This enables us to compute the degree of $\frac{f}{g}$ even when $f$ and $g$ are arbitrary continuous complex functions without common zeros, and the fraction has a limit at infinity (which can be finite or infinite). Specifically, if $f$ and $g$ are complex polynomials in $z$ and $\bar{z}$, and the highest-degree homogeneous component of the polynomial with the greater algebraic degree has a finite or infinite limit as $|z|\to\infty$, then the problem reduces to counting the roots of a complex polynomial inside the unit circle, obtained from this component.
@@ 1. Introduction
In general, there is a standard method for computing the degree of a map. Let $R:S^{n}\to S^{n}$ be a continuous function, and let $p\in S^{n}$ be a regular value. That is, there exists an open neighborhood $U$ of $p$ such that the preimage $R^{-1}(U)$ is a disjoint union of open sets $U_{\alpha}$, and $R$ restricted to each $U_{\alpha}$ is a homeomorphism $U_{\alpha}\to U$. For each $U_{\alpha}$, the restriction $R|_{U_{\alpha}}$ either preserves or reverses orientation. The degree of $R$ is then equal to the number of indices $\alpha$ for which $R|_{U_{\alpha}}$ preserves orientation, minus the number of indices $\beta$ for which $R|_{U_{\beta}}$ reverses orientation, noting that the number of preimages of $p$ is finite. However, this method may be difficult to apply. For a map defined by a rational function in $z$ and $\bar{z}$, finding the preimages of a point may require solving a system of polynomial equations of high degree. Furthermore, computing the sign of the Jacobian determinant at each solution can be computationally prohibitive. For example, let $R:S^{2}\to S^{2}$, where $S^{2}$ is identified with the Riemann sphere, be defined by the rational function
@@
Let $f,g:\mathbb{C}\to\mathbb{C}$ be continuous functions with no common zeros. Assume that $\lim_{|z|\to\infty}\frac{f(z)}{g(z)}=\infty$ and that $f(z)\neq 0$ for all $|z|\geq M$ for some constant $M>0$. Consider the continuous map $R:\bar{\mathbb{C}}\to\bar{\mathbb{C}}$ defined as follows:
@@ 2. Proofs
We identify the sphere $S^{2}$ with the Riemann sphere $\bar{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, via stereographic projection. The sphere can also be viewed as the disk $D^{2}$ with its boundary collapsed to a single point. For convenience, we will write
$R(z)=\frac{f(z)}{g(z)},\qquad R:\overline{\mathbb{C}}\to\overline{\mathbb{C}},$
where $R$ denotes a continuous function, with the understanding that this convention also covers the cases when the denominator vanishes or when $z=\infty$.
@@
$G(z,t)=\begin{cases}\frac{f\left(z\left(\frac{M}{|z|}t+1-t\right)\right)}{|f\left(z\left(\frac{M}{|z|}t+1-t\right)\right)|}|f(z)|,&|z|\geq M\\ f(z),&|z|<M\end{cases}$
It is clear that $|G(z,t)|=|f(z)|$ and $|G(z,t)|\geq M$ for all $t\in[0,1]$. Therefore, the map $G$ is continuous. Moreover, when $|z|\geq M$:
@@
$f_{1}(z)=\begin{cases}f\left(\frac{z}{|z|}M\right),&|z|\geq M\\ f(z),&|z|<M\end{cases}$
@@
Here, the local degree at infinity is $\deg_{\infty}(R)=\text{ind}(f)-\text{ind}(g)$, while for any other preimage $p$, the local degree is $\deg_{p}(R)=\text{ind}_{p}(g)$. Thus, by the argument principle, the sum of local degrees of $g$ is $\text{ind}(g)$, leading to:
$\deg(R)=(\text{ind}(f)-\text{ind}(g))+\text{ind}(g)=\text{ind}(f).$
@@
The function $T$ has no zeros on the unit circle. Suppose, for contradiction, that $T(e^{i\alpha_{0}})=0$ for some $\alpha_{0}\in[0,2\pi]$. Since $T$ is not the zero function, there exists a point $a$ with $T(a)\neq 0$. By the homogeneity of $T$, the limit of $T(z)$ as $|z|\to\infty$ does not exist: along the ray from the origin through $e^{i\alpha_{0}}$ the value of $T$ is identically zero, whereas along the ray through $a$ we have
@@ 3. Examples of Computation
In all these examples, we want to compute the degree of the map $R=\frac{f}{g}$. If $f$ and $g$ are polynomials in $z$ and $\bar{z}$, let $d$ be the degree of the highest-degree polynomial among $f$ and $g$. Let $T(z,\bar{z})$ be the homogeneous polynomial of degree $d$ consisting of the highest-degree homogeneous terms of either $f$ or $g$, depending on which has the greater degree. We then define the polynomial $\tilde{T}(z)=z^{d}T(z,z^{-1})$.
@@
In this example, we consider the polynomial

### v3

@@ Abstract.
We describe an effective method for computing the topological degree of continuous functions $R:S^{2}\to S^{2}$, where $S^{2}$ is the Riemann sphere. Our approach generalizes the degree formula for rational functions of complex polynomials, $\frac{f}{g}$, without common zeros. To apply our method, it is necessary to represent the function $R$ as the ratio of two continuous complex-valued functions $f$ and $g$ without common zeros. By using the Hopf fibration, this method reduces the problem to computing the winding number of a loop. This enables us to compute the degree of $\frac{f}{g}$ even when $f$ and $g$ are arbitrary continuous complex functions without common zeros, and the fraction has a limit at infinity (which can be finite or infinite). Specifically, if $f$ and $g$ are complex polynomials in $z$ and $\bar{z}$, and the highest-degree homogeneous component of the polynomial with the greater algebraic degree has a finite or infinite limit as $|z|\to\infty$, then the problem reduces to counting the roots of a complex polynomial inside the unit circle, obtained from this component.
@@ 1. Introduction
In general, there is a standard method for computing the degree of a map. Let $R:S^{n}\to S^{n}$ be a continuous function, and let $p\in S^{n}$ be a regular value. That is, there exists an open neighborhood $U$ of $p$ such that the preimage $R^{-1}(U)$ is a disjoint union of open sets $U_{\alpha}$, and $R$ restricted to each $U_{\alpha}$ is a homeomorphism $U_{\alpha}\to U$. For each $U_{\alpha}$, the restriction $R|_{U_{\alpha}}$ either preserves or reverses orientation. The degree of $R$ is then equal to the number of indices $\alpha$ for which $R|_{U_{\alpha}}$ preserves orientation, minus the number of indices $\beta$ for which $R|_{U_{\beta}}$ reverses orientation, noting that the number of preimages of $p$ is finite.
However, the standard method may be difficult to apply even in the case $n=2$. For a map between Riemann spheres, defined by a rational function of $z$ and $\bar{z}$, finding the preimages of a point may require solving a system of high-degree polynomial equations. Furthermore, computing the sign of the Jacobian determinant at each solution can be computationally prohibitive. For example, let $R:S^{2}\to S^{2}$, where $S^{2}$ is identified with the Riemann sphere, be defined by the rational function
@@
Let $f,g:\mathbb{C}\to\mathbb{C}$ be continuous functions with no common zeros. Assume that $\lim\limits_{|z|\to\infty}\frac{f(z)}{g(z)}=\infty$ and that $f(z)\neq 0$ for all $|z|\geq M$ for some constant $M>0$. Consider the continuous map $R:\bar{\mathbb{C}}\to\bar{\mathbb{C}}$ defined as follows:
@@ 2. Proofs
We identify the sphere $S^{2}$ with the Riemann sphere $\bar{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, via stereographic projection. The sphere can also be viewed as the disk $D^{2}$ with its boundary collapsed to a single point. For convenience, in the case where a continuous function $R:\overline{\mathbb{C}}\to\overline{\mathbb{C}}$ is given as the ratio of two continuous complex-valued functions $f$ and $g$, we will write
$R(z)=\frac{f(z)}{g(z)},$
with the understanding that this convention also covers the cases when the denominator vanishes or when $z=\infty$.
@@
$G(z,t)=\begin{cases}\frac{f\left(t\frac{Mz}{|z|}+(1-t)z\right)}{|f\left(t\frac{Mz}{|z|}+(1-t)z\right)|}|f(z)|,&|z|\geq M\\ f(z),&|z|<M\end{cases}$
It is clear that $|G(z,t)|=|f(z)|$ for all $t\in[0,1]$. Moreover, if $|z|\geq M$, then for all $t\in[0,1]$ we have
$\left|t\frac{Mz}{|z|}+(1-t)z\right|\geq M.$
Therefore, the map $G$ is continuous. Furthermore, when $|z|\geq M$:
@@
$f_{1}(z)=G(z,1)$
@@
Here, the local degree at infinity is $\deg_{\infty}(R)=\text{ind}_{\infty}(f)-\text{ind}_{\infty}(g)$, while for any other preimage $p$, the local degree is $\deg_{p}(R)=\text{ind}_{p}(g)$. Thus, by the argument principle, the sum of local degrees of $g$ is $\text{ind}_{\infty}(g)$, leading to:
$\deg(R)=(\text{ind}_{\infty}(f)-\text{ind}_{\infty}(g))+\text{ind}_{\infty}(g)=\text{ind}_{\infty}(f).$
@@
Proof.  [of Theorem 2]
The function $T$ has no zeros on the unit circle. Suppose, for contradiction, that $T(e^{i\alpha_{0}})=0$ for some $\alpha_{0}\in[0,2\pi]$. Since $T$ is not identically zero, there exists a point $a$ such that $T(a)\neq 0$. By the homogeneity of $T$, the function $T(z)$ has no limit as $|z|\to\infty$: along the ray from the origin through $e^{i\alpha_{0}}$ the function vanishes identically, whereas along the ray through $a$ we have
Therefore, the minimum of the modulus of $T$ on the unit circle is greater than 0. It then follows from the homogeneity of $T$ that there exists a sufficiently large number $M>0$ such that for all $|z|>M$ the numerator of $R$ has no zeros and
|T(z)|>|f(z)T(z)|.|T(z)|>|f(z)-T(z)|.
@@ 3. Examples of Computation
In all these examples, we want to compute the degree of the map $R=\frac{f}{g}$, where $f$ and $g$ are polynomials in $z$ and $\bar{z}$, with
$d=\max(\deg f,\deg g),$
and define $T(z,\bar{z})$ as the homogeneous polynomial of degree $d$ formed by the highest-degree terms of $f$ or $g$, whichever has the larger degree. We then define $\tilde{T}(z)=z^{d}T(z,z^{-1})$.
@@
The homogeneous polynomial of highest degree is